You 

will combine two numbers from 0 to 9. One entry will win PHP 6000, out of 100 total entries, if 

you match the last two digits of the final score in a basketball game. Let X be the net gain from 

the purchase of one entry. What is the mean or the expected value of a person who buys a ticket? 

Illustrate by constructing a probability distribution. 


1
Expert's answer
2021-05-21T03:13:02-0400

Total number of ways in which we can choose two numbers "= ^{10}C_2 = 45"

One entry will win PHP6000

Probability that will win the entry  "= \dfrac{45}{100} = 0.45"

Probability that does not win "= \dfrac{55}{100} = 0.55"

Probability distribution can be given as,



Expected value of person who buys a ticket "= 0.45 \times 6000 + 0.55 \times 0 = 2700"

Expected value = 2700